Given: AB≅BC ; CD≅BC ;
AB = 2x +1; CD = 4x -11
Prove: x = 6

Answer by ikleyn(52807)

.

     AB≅BC and CD≅BC   implies   AB≅CD.


It means


    2x+1 = 4x-11.


It implies


    1 + 11 = 4x - 2x

    12     = 2x

     x     = 12/2 = 6.


PROVED

Solved, explained, and completed.